Fermi National Laboratory
 

Accelerator Update for May 7, 2007 - May 9, 2007

 

Monday May 7

The day shift began with Operations monitoring Tevatron (TeV) store 5412 (100.95E30) and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI, MiniBooNE, and the Meson MTest experiments taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

At 8:42 AM, a Booster vacuum pump (IPRF17) tripped off and caused two Booster RF stations (BRF7 & 8) to trip off. Everything reset.

At 9:47 AM, Operations reported that cryo system technicians were beginning a TeV sector B0 wet engine repair. D0 and CDF turned off their silicon detectors during the work.

At 10:01 AM, the MI beam permit tripped due to bad beam positions. It reset.

An RF expert reported at 10:02 AM that he had completed work on MIRF14.

At 10:21 AM, MIRF14 tripped off. It reset.

Cryo system techs reported at 10:34 AM that they had completed a bearing replacement on the B0 wet engine.

At 11:11 AM, MIRF14 tripped off. Operations contacted an RF expert who had the station back on line by 12:17 PM.

The Neutron Therapy Facility (NTF) reported at 1:15 PM that they had completed treatment on their last patient for the day.

At 2:29 PM, a Linac RF station (LRF3) tripped off. It reset.

At 3:43 PM, Operations reported having a problem with readbacks for the MI front-end (FE). Reboots didn’t do anything. Operators contacted an expert who fixed the problem.

Operations reported at 4:09 PM that the Recycler had stashed antiprotons twice during the day shift.

The MTest experiment reported at 6 PM that they were finished with the beam for the night.

At 8:55 PM, LRF3 tripped off. It reset. Also, a Booster vacuum pump (IPRF24) tripped off and caused BRF15 & 16 to trip off. BRF16 wouldn’t reset. Operators reduced the booster beam intensity and called in an expert. The expert reset the breaker inside the bias supply. Booster resumed sending its normal beam intensity at 9:50 PM.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Tuesday May 8

The midnight shift began with Operations monitoring store 5412 (34.798E30) and the stash (208.8E10), with Pbar stacking (14.8E10), and with NuMI and MiniBooNE taking beam.

Operations terminated store 5412 at 4:04 AM.

Operations put the TeV into shot setup at 4:16 AM.

Operations established store 5413 at 6:26 AM with an initial luminosity of 181.4E30.

At 7:31 AM, Operations reported that two Recycler transfers were lost due to a power supply (VDPA1) failure. Operators contacted a power supply expert.

Operations reported at 7:56 AM that the Recycler had stashed twice during the midnight shift, but only one stash was successful.

At 8:56 AM, LRF3 tripped off. It reset.

Power supply experts reported that they had retapped the Recycler power supply that had the LCW leak (R:VDPA1). The one good transformer in the supply was tapped to put out 100%. The other transformer is still too wet to use. Experts reported that a replacement supply was being prepared.

A TeV expert began a Tevatron Electron Lens (TEL1) study at 10:27 AM.

At 11:32 AM, LRF2 tripped off. An expert reported that its filament power supply had tripped. He brought it back up.

An MI expert began a barrier bucket study at 1:20 PM.

An MI expert began a parasitic MI-8 collimator study at 2:27 PM.

An MI expert began a slipstacking study at 3:12 PM.

TeV experts began a separator scan study at 4 PM.

Operations reported at 4:08 PM that the Recycler had stashed twice during the day shift.

The TeV separator scans ended at 4:37 PM.

At 5:28 PM, the Linac 750keV area Buncher tripped off. It reset.

MTest reported at 6 PM that they were done with the beam for the night.

At 9:11 PM, LRF3 tripped off. It reset.

At 9:12 PM, KRF2 tripped off due to a LCW flow indication. It reset.

At 9:29 PM, a Recycler CE ( Corrector Element) chassis tripped off. There was no immediate beam loss.

At 9:31 PM, the Recycler had an emittance blow up. There was about 30E10 of antiprotons lost through the duration of the problem. Experts investigated and reported that the blow up was due to the CE chassis trip and specifically due to CE603. They had the chassis back on 30 minutes later.

At 9:34 PM, TEL2 tripped off on a vacuum interlock. It reset.

At 9:57 PM, a bad beam position caused a beam abort at MI-30.

At 11:14 PM, Operations reported that one monitor (LM223) seemed to be causing the MI-30 abort problems. Operators bypassed the monitor at 11:52 PM.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Wednesday May 9

The midnight shift began with Operations monitoring store 5413 (40.58E30) and the stash (168.2E10), with Pbar stacking (34.6E10), and with NuMI and MiniBooNE taking beam.

At 12:48 AM, BRF7 tripped off. It reset.

At 1:01 AM, KRF6 tripped off. It reset.

At 3:44 AM, the Linac Buncher tripped off. It reset.

Operations terminated store 5413 at 4:59 AM.

Operations prepared D0 and CDF for accesses at 5:13 AM.

At 5:54 AM, LRF3 tripped off. It reset.

At 5:55 AM, KRF4 tripped off. It reset.

At 6:20 AM, Operations reported another Recycler emittance blow up. Operators contacted an expert who reported that this blow up was damper related and he was coming in.

MTest began taking beam at 6:29 AM.

A Pbar expert began a reverse proton study at 7:22 AM. There will be no stacking during this study. Operations set up a NuMI only time line.

Recycler experts reported that they reverted to their old damper system at 7:28 AM.

Operations put the TeV through a dry squeeze at 7:29 AM.

Operations reported at 7:55 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Wednesday

The plans for today are to establish a store this morning and to keep it spinning until 5 AM Thursday morning. Pbar will conduct reverse proton studies. Thursday will be full of safety system testing, and with studies and maintenance work scattered around the testing. For more information on the planned schedule of events and for the current situation please check the Run Coordinator’s Electronic Log at:
http://www-bd.fnal.gov/cgi-mach/machlog.pl?nb=runco.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/09/2007    email Fermilab